### v1

# VL-ICL Bench: The Devil in the Details of Benchmarking Multimodal In-Context Learning

Abstract
Large language models (LLMs) famously exhibit emergent in-context learning (ICL) – the ability to rapidly adapt to new tasks using few-shot examples provided as a prompt, without updating the model’s weights. Built on top of LLMs, vision large language models (VLLMs) have advanced significantly in areas such as recognition, reasoning, and grounding. However, investigations into multimodal ICL have predominantly focused on few-shot visual question answering (VQA), and image captioning, which we will show neither exploit the strengths of ICL, nor test its limitations. The broader capabilities and limitations of multimodal ICL remain under-explored. In this study, we introduce a comprehensive benchmark VL-ICL Bench for multimodal in-context learning, encompassing a broad spectrum of tasks that involve both images and text as inputs and outputs, and different types of challenges, from perception to reasoning and long context length. We evaluate the abilities of state-of-the-art VLLMs against this benchmark suite, revealing their diverse strengths and weaknesses, and showing that even the most advanced models, such as GPT-4, find the tasks challenging. By highlighting a range of new ICL tasks, and the associated strengths and limitations of existing models, we hope that our dataset will inspire future work on enhancing the in-context learning capabilities of VLLMs, as well as inspire new applications that leverage VLLM ICL. The code and dataset are available at https://github.com/ys-zong/VL-ICL.
1 Introduction
With the scaling of model size, large language models (LLMs) famously exhibit the emergent capability of in-context learning (ICL) (Brown et al., 2020; Dong et al., 2022). This refers to the ability to learn from analogy within a single feed-forward pass – thus, enabling the model to learn completely new tasks using a few input-output examples, without requiring any updates to the model parameters. This training-free nature of ICL has led to its rapid and broad application across a wide range of scenarios and applications, as illustrated by benchmarks such as (Hendrycks et al., 2021; Zhong et al., 2023; Srivastava et al., 2023; Cobbe et al., 2021).
Vision large language models (VLLMs) are typically built on a base LLM, by augmenting it with a vision encoder and/or decoder connected by some stitching mechanism (Liu et al., 2023a, b; Bai et al., 2023; Li et al., 2023b; Alayrac et al., 2022; Koh et al., 2023; Ge et al., 2024). These models have rapidly advanced alongside LLMs, and attracted significant attention for their remarkable multimodal capabilities in zero-shot recognition, reasoning, grounding, and visual question answering (VQA) among other capabilities. These capabilities have been thoroughly tested by a range of recent benchmark suites (Liu et al., 2023c; Fu et al., 2023; Li et al., 2023c; Yu et al., 2023). Meanwhile, VLLMs are also widely presumed to inherit in-context learning (ICL) capabilities from their base LLM. However, their abilities in this respect are poorly evaluated and poorly understood. Current VLLMs studies mainly report their zero-shot capabilities measured by the benchmarks above, while ICL is usually only evaluated qualitatively, or as a secondary consideration via few-shot visual question answering (VQA) or image captioning (Bai et al., 2023; Awadalla et al., 2023; Sun et al., 2023a; Laurençon et al., 2023), with a notable deficiency in quantitative assessment across a wider spectrum of ICL tasks. This is presumably due to the ready availability of VQA and captioning benchmarking infrastructure. However, we will show that captioning and VQA tasks are not ideal for ICL evaluation: They neither truly exploit the ability of ICL to improve performance from examples; nor they test the limits of what ICL can do, in order to motivate future VLLM research to better exploit and expose the underpinning LLM’s ICL ability.
Figure 1: Illustration of the different tasks in VL-ICL Bench. Image-to-text tasks are in the first three rows, while text-to-image tasks are in the bottom row. Image-to-text tasks in the third row do reasoning on interleaved image-text inputs.
To enhance the understanding of multimodal ICL and assess the ICL capabilities of state-of-the-art VLLMs, we introduce a novel benchmark suite VL-ICL Bench (Figure 1), tailored for assessing VLLM in-context learning. Our benchmark suite incorporates both text-output and image-output tasks, and is designed to test various facets of VLLMs, including fine-grained perception, reasoning, rule induction, and context-length. We conduct comprehensive evaluations of state-of-the-art VLLMs that are capable of processing interleaved image-text as inputs on our benchmark. Results reveal that although certain models exhibit reasonable performance on specific tasks, none demonstrate uniform excellence across the entire spectrum of tasks, and some models perform near chance level on some tasks. We hope that this systematic study of different opportunities and challenges for multimodal ICL will support practitioners to know what is currently possible and impossible in terms of training-free learning of new multimodal tasks, and spur VLLM model developers to study how to expose as much as possible of the LLM’s ICL ability to the multimodal world.
To summarise our contributions: (1) We demonstrate the limitations inherent in the common practice of quantitatively evaluating VLLM ICL via VQA and captioning. (2) We introduce the first thorough and integrated benchmark suite of ICL tasks covering diverse challenges including perception, reasoning, rule-induction, long context-length and text-to-image/image-to-text. (3) We rigorously evaluate a range of state-of-the-art VLLMs on our benchmark suite, and highlight their diverse strengths and weaknesses, as well the varying maturity of solutions to different ICL challenges.
2 Background and Motivation
2.1 The ICL Problem Setting
Given a pre-trained VLLM $\theta$, an optional text instruction $I$, a context set [^1] We use context set and support set interchangeably in this paper. $S=\{(x_{i},y_{i})\}$ of query example $x$ and labels $y$, and a test example $x^{*}$, ICL models estimate
$p_{\theta}(y^{*}|x^{*},I,S)$ (1)
with a feed-forward pass. For LLMs, $x$ and $y$ are typically text. For VLLMs, $x$ can be text and/or images, and $y$ can be text (image-to-text ICL) or images (text-to-image ICL).
This ICL setting is in contrast to the simpler zero-shot scenario, where pre-trained models estimate $p_{\theta}(y^{*}|x^{*},I)$ purely based on the pre-learned knowledge in $\theta$ with no additional training data provided in $S$. The zero-shot scenario has been rigorously evaluated by diverse benchmarks (Liu et al., 2023c; Fu et al., 2023; Li et al., 2023c; Yu et al., 2023), and in the following section we discuss the limitations of existing ICL evaluations that motivate our benchmark.
2.2 Common Practice in ICL Evaluation
The benchmarks that have been most popular in prior attempts at quantitative evaluation of multimodal ICL are VQA and captioning. We focus our discussion in this section on image-to-text models (Alayrac et al., 2022; Bai et al., 2023; Li et al., 2023b; Laurençon et al., 2023; Awadalla et al., 2023), as in-context text-to-image models (Ge et al., 2024; Koh et al., 2023) are relatively less common and less mature, so there is no common evaluation practice yet. In the case of captioning, the context set $S$ contains examples of images $x$ and captions $y$; while for VQA the context $S$ contains image-question pairs $x$ and answers $y$.
Figure 2(a) plots the ICL performance of six state-of-the-art VLLMs on three popular benchmarks – MathVista VQA (Lu et al., 2024), VizWiz VQA (Gurari et al., 2018), and COCO captioning (Lin et al., 2014) for varying numbers of training examples (shots). While the performance of the different models varies, the key observation is that most of the lines shown are only weakly increasing. It means that for all models, there is limited improvement achieved by ICL (shots $>0$) compared to the zero-shot case (shots $=0$). This is because, while the context set $S$ illustrates the notion of asking and answering a question or captioning images, the baseline VLLM $\theta$ is already quite good at VQA and captioning. The limiting factors in VLLM captioning and VQA are aspects such as detailed perception, common sense knowledge, etc. – all of which are fundamental challenges to the VLLM, and not aspects that can reasonably be taught by a few-shot support set.
Given the discussion above, it is unclear why performance should improve with shots at all? We conjectured that this is largely due to the VLLM learning about each dataset’s preferred answer style, rather than learning to better solve multimodal inference tasks per-se. For example, in captioning zero-shot VLLMs tend to produce more verbose captions than the ground truth in COCO, and they learn to be more concise through ICL. Meanwhile, for VQA, there is a standard practice of evaluating based on string match between the ground-truth answer and the model-provided answer. For example, VizWiz has unanswerable questions, which some VLLMs answer with ‘‘I don’t know’’ which would not be a string matched against a ground truth ‘‘Unanswerable’’. Some models thus learn about answer-formatting (e.g., preferred terminology; avoid using any preface or postface that may not satisfy a string match) from the context set. This is indeed a kind of ICL, but perhaps not what one expects to be learning in VQA. To validate this conjecture, we repeat the previous evaluation, but using soft matching to eliminate the impact of answer format learning. For VQA, we use a pretrained LLM to determine whether the prediction is semantically equivalent to the ground-truth while for captioning, we use an LLM to score the quality of the generated caption on a scale of score 1-10 (details in the Appendix). Fig. 2(b) shows that the curves have almost fully flattened out, with zero-shot performance having improved. Fig. 2(c) quantifies this difference by showing the average rate of improvement with shots for exact match and LLM match. The change to LLM validation almost completely eliminates any benefit of ICL over zero-shot.
In contrast to the above, popular LLM ICL benchmarks in the language domain do usually exhibit non-trivial ICL learning (Brown et al., 2020; Dong et al., 2022). Figure 3 shows three state-of-the-art VLLMs along with their corresponding base LLMs, evaluated on three popular NLP tasks (AGNews (Zhang et al., 2015), MIT Movies (Ushio and Camacho-Collados, 2021) and TREC (Voorhees and Tice, 2000)). We can see that in contrast to the VQA/captioning benchmarks, models’ zero-shot performance is often substantially improved by few-shot ICL. This result confirms that the LLM components in VLLMs do inherit the ICL ability of their base LLM. However, it raises the question of how we can meaningfully exploit and measure the ICL ability of VLLMs in the multimodal context. In the next section, we introduce our benchmark VL-ICL Bench, which does exactly this.
(a) Standard metrics for evaluation.
(b) LLM as a judge for evaluation.
(c) Accuracy increments when adding more support examples.
Figure 2: VQA and captioning are poor benchmarks for image-to-text ICL. (a) Evaluating state-of-the-art VLLMs on representative examples of popular image-to-text ICL benchmarks – MathVista, VizWiz, and COCO – with standard evaluation protocol. Zero-shot performance is high, and ICL performance depends only weakly on shots, showing that ICL does not learn much. (b) Re-evaluation of VLLMs with LLM-based evaluation further reduces dependence on shots. (c) The impact of ICL on performance goes from small to negligible when moving from traditional to LLM-based evaluation. ICL on these benchmarks primarily learns answer style/format.
Figure 3: Evaluating state-of-the-art VLLM/LLM pairs on popular text-to-text ICL benchmarks. Few-shot ICL often substantially improves on zero-shot performance, indicating that meaningful in-context-learning is taking place, unlike for the popular image-to-text VLLM benchmarks in Fig. 2.
3 VL-ICL Bench
3.1 Main Multimodal Benchmark
Our VL-ICL Bench covers a number of tasks, which includes diverse ICL capabilities such as concept binding, reasoning or fine-grained perception. It covers both image-to-text and text-to-image generation. Our benchmark includes the following eight tasks: Fast Open MiniImageNet, CLEVR Count Induction, Operator Induction, Interleaved Operator Induction, TextOCR, Matching MiniImageNet, Text-to-image MiniImageNet and CoBSAT. We provide illustrations of the tasks in Figure 1, and summarise the diverse capabilities tested by each VL-ICL Bench task in Table 1. This table also summarises the dataset statistics, demonstrating that VL-ICL Bench is compact and accessible. We follow the typical protocol of the ICL community (Dong et al., 2022; Tsimpoukelli et al., 2021; Min et al., 2022) [^2] This is is different than the few-shot meta-learning community (Wang et al., 2020; Hospedales et al., 2021), which samples support/query sets from the same pool. and split each dataset into train and test splits. Few-shot in-context learning is then performed/evaluated by sampling the support/context set from the training split, and the test/query examples from the testing split. Final performance is the average of a number of such ICL episodes.
Table 1: VL-ICL Bench overview. It evaluates diverse capabilities and challenges of ICL with VLLMs. Meanwhile it is compact and easy to be used by researchers, without prohibitive resource requirements.
Dataset Capabilities Tested Train Set Test Set Size (GB)
Fast Open MiniImageNet I2T, Fast Binding 5,000 200 0.18
CLEVR Count Induction I2T, Fine Grained Perception, Induction 0,800 200 0.18
Operator Induction I2T, Induction, Reasoning 00,80 060 0.01
Interleaved Operator Induction I2T, Induction, Reasoning, Interleaving, Long-Context 00,80 060 0.01
TextOCR I2T, Fine Grained Perception, Induction 0,800 200 0.98
Matching MiniImageNet I2T, Induction, Interleaving, Long-Context 1,600 400 0.11
Text-to-image MiniImageNet T2I, Fast Binding 5,000 200 0.18
CoBSAT T2I, Induction 0,800 200 0.07
Total T2I, I2T, Binding, Perception, Long-Context, Interleaving, Induction, Reasoning 14,160 1,520 1.72
Fast Open MiniImageNet  We use the variant of MiniImageNet few-shot object recognition (Vinyals et al., 2016) repurposed for ICL in (Tsimpoukelli et al., 2021). In open-ended classification VLLMs must learn to open-endedly name an object based on a few examples, rather than simply classifying it into a pre-defined set of options. Thus the chance rate is effectively zero, rather than dependent on the number of categories. Fast-binding tasks test the ability of models to associate novel names or symbols to concepts, without relying on prior knowledge. Thus, (Tsimpoukelli et al., 2021) give synthetic names to object categories in the context/support set (e.g. dax or perpo), and the model must learn to associate the names with the illustrated visual concept in order to correctly name the test images. We use the two-way version.
CLEVR Count Induction  In this dataset, models must learn to solve tasks of the type ‘‘How many red objects are there in the scene?’’. However, they must learn this from example rather than being explicitly prompted to do so. Specifically, we input $x$ scene images from CLEVR (Johnson et al., 2017), along with an attribute: value pair that identifies a particular type of object within the scene. There are four types of attributes: size, shape, color, material. The required output $y$ is the count of the objects in the image that meet the provided attribute: value criterion. To solve this task, the model has to provide fine-grained perception to support counting and differentiating object types, learn to ground the requested attribute in images, and importantly induce that the required operator mapping $x$ to $y$ is counting the specified object type [^3] This task could be performed zero-shot with a suitably detailed VQA prompt. However, the goal is to test whether models can learn the task from a few examples by ICL..
Operator Induction  In this image-to-text task, models must solve tasks of the type 2 ? 7 = 9 given training examples like 1 ? 3 = 4. It means that besides parsing an image $x$ to extract the numbers and the operator, models need to induce that the role of the unknown operator is addition, and then conduct simple arithmetic reasoning to apply the induced operator on the parsed test examples. Available mathematical operations are plus, minus and times, and we only consider single digit numbers. We generate our own images for this task.
Interleaved Operator Induction  This task tests the ability of models to reason over multiple images within $x$ to produce a single answer $y$. In this variation of operator induction we give the model two query images as input containing each number in the expression, rather than a single image containing the whole expression, as above. Otherwise it is the same as the basic operator induction task. In one sense, separating the images makes the task easier, as it substantially simplifies the perception task of parsing the expression from a single image. However, it is also harder in that it requires the VLLM to perform induction and reasoning between two different images, when VLLM training (e.g., captioning) typically trains models to produce outputs that depend on a single image at a time. By introducing multiple images, it also introduces a greater number of tokens than basic operator induction, and thus stresses VLLMs’ ability to use a larger context length.
TextOCR  We repurpose the TextOCR dataset (Singh et al., 2021) to create a task where the model should learn to output the text shown in the red rectangle, as inspired by (Sun et al., 2023a). Images $x$ contain an image with a window of text highlighted, and outputs $y$ are the OCR text. This task could be achieved by a suitably detailed zero-shot prompt, but unlike (Sun et al., 2023a) we focus on evaluating whether the task can be induced by way of example through ICL. Thus this task tests both fine-grained perception and induction capabilities.
Matching MiniImageNet  This task is the simplest example of supervised learning of a relation between two images. For relation learning, inputs contain an image pair $x=\{x_{1},x_{2}\}$, and output $y$ indicates whether a specific relation $r$ holds between them. VLLMs are required to learn the relation from examples where it does ($r(x_{1},x_{2})=true$) and does not ($r(x_{1},x_{2})=false$) hold. We re-use MiniImageNet (Tsimpoukelli et al., 2021; Vinyals et al., 2016) dataset and the relation to learn is whether the two images come from the same class or not (Sung et al., 2018), with disjoint sets of classes considered between training and testing. This task tests induction, the ability to process multiple interleaved images, and the ability to process larger context lengths.
Text-to-Image MiniImageNet  This novel task tests fast concept binding in the text-to-image context. We introduce a further variation of MiniImageNet (Tsimpoukelli et al., 2021; Vinyals et al., 2016), which inputs synthetic category names $x$ (for fast binding), and outputs images $y$. The model should learn from the context set to associate synthetic names with distributions over images, and thus learn to generate a new image of the corresponding category when prompted with the artificial category name (Figure 1). This task tests image generation and fast binding. LLaVA-Next-7B is used to assess generation correctness.
CoBSAT  We finally also utilize a recent text-to-image CoBSAT (Zeng et al., 2024) benchmark as part of our larger VL-ICL Bench suite (Figure 1). This is a text-to-image task where the model must learn to synthesise images $y$ of a specified text concept $x$ (e.g., object category), but furthermore there is a latent variable common to the context set examples that must be induced and correctly rendered in novel testing images (e.g., common color of objects). This task tests image generation and latent variable induction.
Capability Summary  The VL-ICL Bench suite described above goes far beyond any individual existing ICL benchmark to test diverse capabilities of multimodal ICL including (Table 1): Both text-to-image and image-to-text generation; fast-binding – the ability to rapidly ground new symbols to visual concepts and re-use those symbols in the context of new data; fine-grained perception – as required to count or read text; interleaving – the ability to reason over the content of multiple images when generating a single output; rule induction – inducing non-trivial concepts such as mathematical operators and latent variables from examples; simple reasoning such as arithmetic; and long-context – the ability of a VLLM to usefully exploit a large number of context tokens.
3.2 Text Variation
In order to compare the impact of multimodality, we also include text-version alternatives for our tasks. For datasets such as open-ended MiniImageNet, instead of images we provide image captions and use those for reasoning. For example, in CLEVR we provide enumerations of the objects in the scene, including their attributes. Note that text versions are not practical for all of the tasks, in particular, TextOCR is difficult to translate into a suitable text alternative.
4 Results
4.1 Experiment Setup
Models  We evaluate a diverse family of state-of-the-art models with various sizes (ranging from 7B to 80B) and different LLM backbones on our benchmark. Specifically, for image-to-text VLLMs, we select Open Flamingo (9B) (Awadalla et al., 2023), IDEFICS (9B and 80B) (Laurençon et al., 2023), Otter (9B) (Li et al., 2023b), InterLM-XComposer2 (7B) (Zhang et al., 2023), LLaVA-Next (Vicuna-7B) (Liu et al., 2024a), Qwen-VL-Chat (9B) (Bai et al., 2023), and Emu2-Chat (34B) (Sun et al., 2023a). For text-to-image VLLMs, we use GILL (7B) (Koh et al., 2023), SEED-LLaMA (8B, 14B) (Ge et al., 2024), Emu1 (14B) (Sun et al., 2023b), Emu2-Gen (34B) (Sun et al., 2023a). We also evaluate GPT4V (OpenAI, 2023) on our benchmark. We use officially released model weights or GPT4 API and adopt greedy decoding for reproducibility. All experiments are conducted using three different random seeds and we report the average performance. A100-80GB GPUs are used for experiments.
Prompt  For consistency, we employ the following prompt format for in-context learning. Additionally, we investigate the impact of various prompt formats, with detailed findings presented in the supplementary materials.
[Task Description] Support Set: [Image][Question][Answer] (n-shot) Query: [Image][Question] Prediction: [Answer]
Figure 4: VL-ICL Bench results. Top two rows: Image-to-Text. Bottom: Text-to-Image tasks.
4.2 Main Results
Figure 5: Illustration of how the best models perform on our benchmark. Left: By dataset separately. Right: By capability evaluated, averaging over datasets.
The main results for VL-ICL Bench are presented in Fig. 4 including a breakdown over shots, and summarised for the 2-shot case as a radar plot over tasks and capabilities in Fig. 5. We make the following observations: (1) VLLMs demonstrate non-trivial in-context learning on VL-ICL Bench tasks. Unlike the common VQA and captioning benchmarks (Fig. 2), our tasks have low zero-shot performance and in every task at least one model shows a clear improvement in performance with number of shots. Thus, ICL capability is now indeed being demonstrated and exploited. (2) VLLMs often struggle to make use of a larger number of ICL examples. For several tasks and models performance increases with the first few shots; but the increase is not monotonic. Performance often decreases again as we move to a larger number of shots (e.g., GPT4V CLEVR Count Induction; InternLM-XComposer2 Operator induction; IDEFICS-80B Interleaved operator induction). Models are eventually confused by this greater number of images and tokens, rather than exploiting them to learn the task at hand. This is exacerbated by the difficulty of extrapolation over context length and number of input images, which for higher-shot ICL becomes greater than the context length and image number used for VLLM training. This shows an important limit of the current state-of-the-art in ICL: Future models must support longer contexts and more images to benefit from larger support sets. (3) GPT4V is the best overall image-to-text model. Among all of the models GPT4V is the strongest all round (but surpassed by some such as OpenFlamingo in low-shot MiniImageNet). (4) Zero-shot performance is not strongly indicative of ICL ability. LLaVA-Next-7B (Liu et al., 2024a) is perhaps worst overall on VL-ICL Bench, which is surprising as it is a state-of-the-art open-source model in mainstream zero-shot benchmarks. This is due to point (2): Its training protocol uses one image at a time, and it uses a large number of tokens per image – thus ICL requires it to extrapolate substantially in input image number and token number, which it fails to do. (5) There is No clear winner among text-to-image models. However, text-to-image models have more consistent shot scaling than image-to-text models. This is due to training with more diverse interleaved datasets that provide multiple input images per instance, and using fewer tokens per image for better scaling.
4.3 Additional Analysis
We next use VL-ICL Bench to analyse the role of several challenges and factors influencing ICL performance.
Fast Concept Binding  In our open miniImageNet task, we follow (Tsimpoukelli et al., 2021) to require fast-binding of synthetic concept names so as to purely test models’ ICL ability, without confounding by VLLMs’ zero-shot ability to associate visual concepts with names. Fig. 6 compares the fast and real-world miniImageNet recognition, where we see the fast-binding case is more challenging.
Figure 6: Comparison of fast binding vs real-names version of MiniImageNet. Fast-binding has zero accuracy for zero-shot inference, unlike the real-names version. It is thus solely dependent on ICL for success.
Direct Comparison of Multimodal and Text ICL  We can disentangle the role of text versus image inputs for some image-to-text VL-ICL Bench tasks, where we can easily provide a semantically equivalent text input describing the image, in place of image tokens. Fig. 7 shows a comparison between image-input vs text-input for count induction, operator induction, and interleaved operator induction tasks. With text input, performance grows much more sharply and consistently with the number of shots. This is attributable to both (i) reduction of perception difficulty, and (ii) reduction in the total number of tokens compared to image input.
Figure 7: Comparison of multimodal (dashed line) and text (solid line) ICL. Performance increases more sharply and consistently with text inputs.
Scaling with Number of Shots  As discussed in Sec. 4.2, the various models exhibit different scaling abilities with respect to number of shots. We summarise their scaling ability by aggregating over tasks and reporting the average accuracy increment in each shot increment Fig. 8. Evidently, VLLMs vary in their accuracy increment per shot, and how well they can extract knowledge from a growing number of shots.
Figure 8: Aggregate analysis across datasets to find the average performance increments when more shots are added. The individual bars of one color correspond to improvements from 0$\rightarrow$1, 1$\rightarrow$2 and 2$\rightarrow$4.
4.4 Qualitative Analysis
We also include a qualitative analysis, where we analyse the impact of using more support examples on the quality of the output. We analyse text-to-image tasks in Fig. 9, using Emu2-Gen model. For text-to-image MiniImageNet the model should learn from the support examples that the artificial names slation and shously correspond to a lion and a school bus, respectively. Emu2-Gen is able to do it to a certain extent, but may get confused by additional support examples as more support examples are not necessarily helpful. In CoBSAT, the support set induces that the animal should have a glacier and desert background. With no support examples the model only displays the animal, but with more support examples it learns that it should use glacier background in the first example. In the second example, the model is able to capture that it should use desert background, but is less successful in showing the required animal – zebra. The quality of the generated images is not necessarily better with more support examples.
Figure 9: Qualitative analysis: Images generated by Emu2-Gen show the ability to learn the concept induced by the support examples.
For qualitative analysis of image-to-text tasks, we discuss some of the common mistakes that the models make for each task.
Open-Ended MiniImageNet  It is relatively common for the models to predict the real-world class, even if it is asked to use the artificial names from the support set. With more support examples such mistakes are less likely to occur as the model learns to use the artificial names.
CLEVR Count Induction  In many cases the model rephrases the question, while in others it says that e.g. the described object is present. Such behaviour is more common with fewer or no support examples. With more support examples the model learns to predict a count but gets incorrect answer. It can be because some objects are more difficult to recognize, e.g. if one partially covers another.
Operator Induction  A very common mistake is to use a different operator than what would be induced from the support examples. For example, the model may guess it should add two numbers instead of multiplying them and vice versa.
Interleaved Operator Induction  The model sometimes predicts the first displayed number or a direct combination of them, e.g. if the two numbers are 1 and 2, it returns 12. It is also relatively common to use an incorrect operator between the numbers.
TextOCR  In many cases the model returns more words than are highlighted in the red box, but includes the highlighted word as one of them. It is also common that the model misses a letter in the text or returns a word that is similar but different from the correct answer. In some cases though the answer may be very different from what is highlighted, possibly returning a different word in the image.
Matching MiniImageNet  The models often describe what is shown on one of the images. However, in many cases they simply return the wrong answer, saying no when it should be yes.
5 Related Work
VLLM Evaluation  With the rapid development of VLLMs, researchers are creating evaluation benchmarks to thoroughly assess the capabilities of VLLMs from diverse perspectives. These evaluations range from zero-shot aggregated benchmarks such as MME (Fu et al., 2023), MMbench (Liu et al., 2023c), and MM-VET (Yu et al., 2023) to datasets designed for fine-tuning on specific aspects, such as visual reasoning (Hudson and Manning, 2019) and knowledge-grounded QA (Lu et al., 2022). They predominantly focus on single-image scenarios, leaving in-context learning evaluation underexplored.
In-Context Learning Evaluation  The term “in-context” has been used in a few ways, including to describe scenarios with interleaved inputs, such as multiple video frames or multi-turn conversations (Li et al., 2023a, b; Zhao et al., 2023; Ge et al., 2024). Although the study of interleaved inputs presents an intriguing subject, it does not align with the core definition of in-context learning that we consider following (Brown et al., 2020; Dong et al., 2022; Min et al., 2022), which involves the emergent ability to learn a function from $x\to y$ from few-shot support input-output pairs. Prior evaluation of ICL (Awadalla et al., 2023; Bai et al., 2023; Laurençon et al., 2023; Sun et al., 2023a) in this sense is limited, and comes with serious drawbacks as discussed in Sec. 2.2. Concurrent to our work, CobSAT (Zeng et al., 2024) introduces a benchmark designed to evaluate in-context learning in text-to-image models, focusing particularly on latent variable induction capabilities. Our work expands upon this by encompassing tasks for both image-to-text and text-to-image generation, assessing a wider array of capabilities (Tab. 1). Additionally, we have incorporated CobSAT as a subset of our benchmark.
Visual In-Context Learning  The term “in-context” has also been used in pure vision models, which aim to perform diverse image-to-image tasks without task-specific prediction heads (Bar et al., 2022; Wang et al., 2023a, b), such as semantic segmentation, depth estimation, object detection, etc. However, these models are explicitly trained on paired in-context input-output data to be able to perform visual ICL during inference. In this paper, we focus on multimodal vision-language ICL, which is based on the emergent ability of LLMs.
6 Conclusion
We have introduced the first comprehensive benchmark suite VL-ICL Bench for multimodal vision-and-language in-context learning with VLLMs. This benchmark suite avoids the issue with the existing mainstream but limited approach to evaluating image-to-text ICL – that ICL provides limited demonstrable benefit over zero-shot inference, and VLLMs learn answer formatting at best rather than any true multimodal capability. In contrast, VL-ICL Bench tests a wide variety of multimodal capabilities including both text-to-image and image-to-text generation, fine-grained perception, rule-induction, reasoning, image interleaving, fast concept binding, long context, and shot scaling. We hope this benchmark will inspire model developers to consider all these capabilities in VLLM development, and inform practitioners about the evolution of what VLLM ICL can and cannot do as the field develops. We also aim to expand our benchmark to incorporate more tasks and models in the future.
Appendix
  • Section 1: Implementation and Evaluation Details
  • Section 2: Further Experimental Analysis
    Subsection 2.1: Results of Scaling to Many Shots (Figure 10).
    Subsection 2.2: Results of Chain-of-Thought Prompting (Figure 11).
    Subsection 2.3: Results of Repeating the Support Set (Figure 12).
    Subsection 2.4: Influence of Instruction Fine-tuning to ICL (Figure 13).
    Subsection 2.5: Different Levels of Task Descriptions (Figure 14).
  • Section 3: Complete Results
    Raw results of all figures in the main text and appendix from Table 3 to 50.
1 Implementation and Evaluation Details
VL-ICL Bench Evaluation Metrics  We use accuracy as the metric across all subsets in our benchmark. For text-to-image generation tasks, we utilize the state-of-the-art VLLM LLaVA-Next [Liu et al., 2024a] as the judge model to decide whether the generated images contain the required object or attribute.
Models Configurations  We additionally provide a summary of the configurations for the models benchmarked in our paper in Table 2, with a particular focus on the number of tokens per image and the context length. This information helps elucidate why some models exhibit poor scalability with increasing shots, as the total lengths exceed the maximum context window.
Table 2: Detailed configurations of the models used in our benchmark.
Model Connection Module Image Tokens Context Length (Train) Context Length (Test)
OpenFlamingo-9B Perceiver 64 2048 2048
IDEFICS-9B Perceiver 64 2048 2048
Otter Perceiver 64 2048 2048
InternLM-XComposer2 Perceiver 64 2048 4096
Qwen-VL-Chat Cross-Attention 256 2048 8192
LLaVA-Next MLP 576 2048 4096
Emu1 C-Former 512 2048 2048
Emu2 Linear layers 64 2048 2048
GILL Linear layers 4 2048 2048
SEED-LLaMA Q-Former 32 2048 4096
Prompts  We list specific prompts below that we use for specific experiments.
Prompt to judge image generation for Fast MiniImageNet and CobSAT dataset User: Decide whether the image contains the following concept: {GT}. Answer with ’yes’ or ’no’.
Prompt to judge the answer for Vizwiz VQA. User: Based on the image and question, decide whether the predicted answer has the same meaning as the ground truth. Answer with ’yes’ or ’no’. Question: {Question} Predicted answer: {Prediction} Ground Truth: {GT}
Prompt to rate the quality of COCO captioning (Main text, section 2.2) User: Given the following image, you are to evaluate the provided generated caption based on its relevance, accuracy, completeness, and creativity in describing the image. Rate the caption on a scale from 1 to 10, where 10 represents an exceptional description that accurately and completely reflects the image’s content, and 1 represents a poor description that does not accurately describe the image. Generated Caption: {Prediction} Ground Truth Caption: {GT} Consider the following criteria for your rating: 1 (Very Poor): The caption does not correspond to the image’s content, providing incorrect information or irrelevant descriptions. It misses essential elements and may introduce non-existent aspects. 3 (Poor): The caption only slightly relates to the image, missing significant details or containing inaccuracies. It acknowledges some elements of the image but overlooks key aspects. 5 (Fair): The caption provides a basic description of the image but lacks depth and detail. It captures main elements but misses subtleties and may lack creativity or precision. 7 (Good): The caption accurately describes the main elements of the image, with some attention to detail and creativity. Minor inaccuracies or omissions may be present, but the overall description is sound. 8 (Very Good): The caption provides a detailed and accurate description of the image, with good creativity and insight. It captures both essential and minor elements, offering a well-rounded depiction. 9 (Excellent): The caption delivers an accurate, detailed, and insightful description, demonstrating high creativity and a deep understanding of the image. It covers all relevant details, enhancing the viewer’s perception. 10 (Exceptional): The caption offers a flawless description, providing comprehensive, accurate, and highly creative insights. It perfectly aligns with the image’s content, capturing nuances and offering an enhanced perspective. Please provide your rating. You should ONLY output the score number.
2 Further Analysis
2.1 Scaling to More Shots
To examine the maximum number of shots the models can handle and whether the model can still benefit from more shots, we further increase the support set size to 16, 32, and 64 shots. We choose three models for this experiment: OpenFlamingo 9B [Awadalla et al., 2023], IDEFICS-9B-Instruct [Laurençon et al., 2023], and InternLM-XComposer2 [Zhang et al., 2023]. These models were selected because each image they process translates to fewer tokens (Table 2), ensuring that they do not exceed the maximum context length when evaluated with 64 shots. IDEFICS-9B-Instruct demonstrates a better scaling capability compared to other models in most of the datasets. Besides, while InternLM-XComposer2 has strong performance in a low-shot regime, the performance quickly decreases with many shots. This may be due to the mismatch between training (4096) and testing (2048) context length (Table 2) where the extrapolation of context length has been a well-known challenging task [Press et al., 2022, Liu et al., 2024b].
Figure 10: Results of scaling to many shots (Max 64). IDEFICS-9B-Instruct exhibits strong scaling capabilities across most datasets compared to other models. Additionally, while InternLM-XComposer2 shows strong performance in low-shot scenarios, its performance diminishes rapidly as the number of shots increases.
2.2 Chain-of-Thought Prompting
To investigate whether there is any strategy that can enhance in-context learning, one straightforward method is Chain-of-Thought (CoT) prompting [Wei et al., 2022]. CoT prompts the model to articulate its reasoning process concerning latent variables from the support set, potentially improving its learning and inference capabilities. We experiment with Qwen-VL-Chat [Bai et al., 2023] and InternLM-XComposer2 [Zhang et al., 2023] that have state-of-the-art LLMs with strong reasoning ability. Below is the specific prompt we use. [CoT Prompt]: Let’s first think step by step and analyze the relationship between the given few-shot question-answer pairs. Give reasoning rationales. User: [Task Description][Support Set][Query][CoT Prompt] VLLMs: [Generated rationals] User: [Task Description][Support Set][Query][Generated rationals] VLLMs: Prediction
We do not observe a consistent improvement with chain-of-thought prompting: it benefits performance on some datasets while detracting from it on others. These findings underscore the complexity of in-context learning tasks, suggesting that fundamental advancements in model development are necessary. Such tasks cannot be readily addressed with simple prompting techniques like CoT.
Figure 11: Comparison of Chain-of-Thought prompting (dashed line, diamond markers) with baseline results (solid line, circle markers) across a selection of datasets and models. Chain-of-thought prompting does not consistently improve performance across datasets, highlighting the complexity of in-context learning tasks and the need for fundamental model development beyond simple prompting techniques.
2.3 Repeating Support Set
In this subsection, we experiment with an interesting setting: we duplicate the same support example multiple times to assess whether repetition enhances performance. We employ the Qwen-VL-Chat model for these experiments, with the results presented in Figure 12. We found that duplicating shots is particularly beneficial in the 1-shot scenario for Fast Open-Ended MiniImageNet, although this is not consistently observed across other datasets. The likely reason is that Fast Open-Ended MiniImageNet gains from the reinforcement of binding the concept through repeated examples, whereas for tasks like operator induction, diverse examples are necessary to facilitate the learning process.
Figure 12: Investigation of the impact of repeating the in-context examples across a selection of datasets, using Qwen-VL-Chat model. The X-axis represents the number of unique shots, not the total number of shots. For example, 1-shot Repeat x2 means there is one unique shot and it is repeated twice.
2.4 Influence of Instruction Fine-tuning
We investigate how instruction-following fine-tuning affects in-context learning capabilities. We compare two model families, each with a pre-trained version and an instruction-following fine-tuned version: Qwen-VL versus Qwen-VL-Chat [Bai et al., 2023] and IDEFICS-9B versus IDEFICS-9B-Instruct [Laurençon et al., 2023]. Their performance differences are illustrated in Figure 13. Although the outcomes vary, models not fine-tuned with instructions exhibit marginally better scalability concerning the number of shots, as seen with the TextOCR dataset. Further studies are needed to understand whether instruction-following fine-tuning harm the in-context ability.
Figure 13: Comparison of using (solid line) and not using instruction tuning (dashed line). Although the outcomes vary, models not fine-tuned with instructions exhibit marginally better scalability with respect to the number of shots, as evidenced in datasets like TextOCR.
2.5 Different Levels of Task Description Details
We show the impact of different levels of details in the prompt description in Figure 14. The results show that generally the best results are obtained with the most detailed descriptions, but this is not necessarily the case in all settings and in some cases, even no descriptions can be better. The performance is often similar across different levels of details, but in some cases, it can be significantly worse, e.g. for TextOCR. We also provide tables with the full results. In our main experiments, we adopt detailed task descriptions for all datasets.
The task descriptions that we use for the different datasets are as follows:
Fast Open-Ended MiniImageNet Detailed: Induce the concept from the in-context examples. Answer the question with a single word or phase. Concise: Answer the question with a single word or phase.
CLEVR Count Induction Detailed: The image contains objects of different shapes, colors, sizes and materials. The question describes the attribute and its value. You need to find all objects within the image that satisfy the condition. You should induce what operation to use according to the results of the in-context examples and then calculate the result. Concise: Find objects of the given type, induce what operation to use and calculate the result.
Operator Induction Detailed: The image contains two digit numbers and a ? representing the mathematical operator. Induce the mathematical operator (addition, multiplication, minus) according to the results of the in-context examples and calculate the result. Concise: Induce the mathematical operator and calculate the result.
TextOCR Detailed: An image will be provided where a red box is drawn around the text of interest. Answer with the text inside the red box. Ensure that the transcription is precise, reflecting the exact characters, including letters, numbers, symbols. Concise: Answer with the text inside the red box.
Figure 14: Comparison of detailed task description (solid line, circle markers), concise task description (dashed line, x markers) and no task description (dotted line, diamond markers) across a selection of datasets and models.
3 Complete Results
In this section, we show the raw results of the figures in the main text in Section 3.1 to 3.3, and results of supplementary materials in Section 3.4.
3.1 Initial Analysis of VQA and Image Captioning
Table 3: Results of MathVista using string match.
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
OpenFlamingo-9B 9.00 17.00 20.00 21.80 22.00
Otter 14.00 16.50 18.00 18.50 19.50
IDEFICS-9B 0.50 5.50 16.60 20.50 21.00
InternLM-XComposer2 22.50 21.00 22.50 24.00 24.00
Qwen-VL-Chat 20.00 21.50 22.00 22.00 22.00
LLaVA-Next-Vicuna-7B 23.00 22.00 15.00 10.00 9.50
Emu2-Chat 18.50 22.50 23.50 23.50 20.00
Table 4: Results of MathVista using LLM for answer extraction.
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
OpenFlamingo-9B 21.00 22.00 22.00 22.00 23.50
Otter 20.00 20.00 17.50 19.50 19.50
IDEFICS-9B 22.50 22.00 22.00 23.50 23.50
InternLM-XComposer2 28.00 28.50 28.50 29.50 26.00
Qwen-VL-Chat 19.00 21.00 19.50 18.50 19.00
LLaVA-Next-Vicuna-7B 23.00 25.00 18.00 15.00 10.50
Emu2-Chat 24.50 22.50 23.50 23.50 22.00
Table 5: Results of VizWiz using exact match.
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
OpenFlamingo-9B 24.57 29.31 30.62 32.14 35.11
Otter 20.13 24.21 25.78 26.33 25.71
IDEFICS-9B 33.20 31.67 45.33 42.80 41.87
InternLM-X2 71.93 66.33 66.73 66.27 71.73
Qwen-VL-Chat 32.40 34.80 35.20 40.80 40.90
LLaVA-Next-Vicuna-7B 54.12 28.13 10.20 6.60 0.40
Emu2-Chat 31.06 34.20 36.13 40.12 42.66
Table 6: Results of VizWiz using LLM as the judge.
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
OpenFlamingo-9B 31.54 35.85 36.10 36.4 37.28
Otter 23.40 24.12 25.88 26.19 24.93
IDEFICS-9B 35.10 36.90 40.66 42.30 41.35
InternLM-XComposer2 72.90 68.20 69.10 66.30 68.42
Qwen-VL-Chat 37.40 37.53 39.22 40.38 41.20
LLaVA-Next-Vicuna-7B 55.26 26.08 11.38 8.72 2.50
Emu2-Chat 35.68 35.83 37.67 40.51 41.99
Table 7: Results of COCO captions (CIDEr).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
OpenFlamingo-9B 78.2 85.6 88.1 89.0 93.3
Otter 86.2 91.0 97.0 98.6 100.5
IDEFICS-9B 71.6 84.20 90.8 93.1 98.5
InternLM-XComposer2 125.74 125.26 138.82 135.22 129.8
Qwen-VL-Chat 121.10 132.6 135.3 135.4 136.9
LLaVA-Next-Vicuna-7B 131.24 81.75 40.49 34.47 26.26
Emu2-Chat 126.3 127.0 132.06 131.10 133.5
Table 8: Results of COCO captions. Scores are rated by LLaVA-Next from 1-10 (higher the better).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
OpenFlamingo-9B 2.71 2.42 2.75 2.88 2.90
Otter 5.60 5.78 5.68 5.72 5.56
IDEFICS-9B 6.15 6.22 6.15 6.21 6.22
InternLM-XComposer2 6.67 4.17 4.67 5.11 4.58
Qwen-VL-Chat 6.17 6.17 6.12 6.11 6.12
LLaVA-Next-Vicuna-7B 6.34 3.54 3.75 3.43 3.57
Emu2-Chat 6.15 5.89 6.05 6.11 6.15
Table 9: Comparisons of VLLMs and LLMs for text ICL on AGNews dataset (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
LLaVA-Next-Vicuna-7B 74.66 81.16 81.61 84.05 85.38
Vicuna-7B 65.83 79.22 80.20 82.24 82.41
Qwen-VL-Chat 75.49 72.74 73.78 78.62 80.91
QwenLM-7B 54.80 59.51 67.53 72.80 74.64
InternLM-XComposer2 83.99 82.87 82.80 83.28 83.97
InternLM2-Chat-7B 81.89 84.11 84.25 84.32 84.33
Table 10: Comparisons of VLLMs and LLMs for text ICL on MIT Movies dataset (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
LLaVA-Next-Vicuna-7B 47.47 29.88 59.04 85.06 89.16
Vicuna-7B 50.36 63.61 77.83 86.99 89.88
Qwen-VL-Chat 50.36 26.99 46.02 74.70 85.54
QwenLM-7B 66.27 46.99 64.10 85.30 92.53
InternLM-XComposer2 69.64 69.64 78.31 88.19 90.12
InternLM2-Chat-7B 61.45 63.61 73.98 87.71 94.46
Table 11: Comparisons of VLLMs and LLMs for text ICL on TREC dataset (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
LLaVA-Next-Vicuna-7B 49.80 51.80 55.60 65.40 71.00
Vicuna-7B 46.20 55.00 63.60 64.20 70.40
Qwen-VL-Chat 43.60 51.00 55.40 61.20 72.60
QwenLM-7B 40.80 54.00 61.00 63.00 73.90
InternLM-XComposer2 59.00 71.20 75.00 81.80 82.80
InternLM2-Chat-7B 62.00 80.20 85.00 87.00 87.20
3.2 Main Results
We present the main results from Table 12 to 24.
Table 12: Results of different models on Fast Open-Ended Mini-ImageNet (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 5-Shot
OpenFlamingo-9B 0.00 $\pm$ 0.00 39.50 $\pm$ 1.22 58.17 $\pm$ 3.57 51.17 $\pm$ 0.85 54.50 $\pm$ 5.66
IDEFICS-9B 0.00 $\pm$ 0.00 22.00 $\pm$ 0.41 52.00 $\pm$ 2.94 53.83 $\pm$ 0.94 59.17 $\pm$ 6.20
IDEFICS-80B 0.00 $\pm$ 0.00 28.50 $\pm$ 0.27 49.50 $\pm$ 1.28 52.47 $\pm$ 3.25 62.50 $\pm$ 2.00
Otter 0.00 $\pm$ 0.00 10.00 $\pm$ 0.71 25.00 $\pm$ 1.22 28.50 $\pm$ 2.86 25.67 $\pm$ 2.25
InternLM-X2 0.00 $\pm$ 0.00 14.83 $\pm$ 1.03 38.00 $\pm$ 1.78 49.00 $\pm$ 1.78 50.33 $\pm$ 3.86
Qwen-VL-Chat 0.00 $\pm$ 0.00 0.50 $\pm$ 0.41 47.33 $\pm$ 2.49 58.00 $\pm$ 2.83 55.17 $\pm$ 2.25
LLaVA-Next-7B 0.00 $\pm$ 0.00 22.17 $\pm$ 4.03 33.67 $\pm$ 2.25 0.00 $\pm$ 0.00 0.33 $\pm$ 0.24
Emu2-Chat 0.00 $\pm$ 0.00 8.00 $\pm$ 1.87 29.33 $\pm$ 1.84 28.18 $\pm$ 4.26 27.54$\pm$ 5.12
GPT4V 0.00 14.00 48.00 56.00 78.00
Table 13: Results of different models on Real-name Mini-ImageNet (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 5-Shot
OpenFlamingo-9B 0.00 $\pm$ 0.00 26.00 $\pm$ 2.86 53.33 $\pm$ 3.27 52.83 $\pm$ 0.94 49.50 $\pm$ 1.22
IDEFICS-9B 26.50 $\pm$ 0.00 41.83 $\pm$ 2.25 74.50 $\pm$ 2.27 89.00 $\pm$ 0.41 91.17 $\pm$ 1.89
IDEFICS-80B 30.50 $\pm$ 0.00 41.83 $\pm$ 1.18 82.00 $\pm$ 2.68 94.67 $\pm$ 0.62 91.33 $\pm$ 1.43
Otter 13.00 $\pm$ 0.00 51.00 $\pm$ 2.16 57.33 $\pm$ 3.09 56.50 $\pm$ 1.08 61.00 $\pm$ 1.87
InternLM-X2 20.00 $\pm$ 0.00 31.50 $\pm$ 1.63 67.00 $\pm$ 1.47 66.83 $\pm$ 0.24 66.67 $\pm$ 1.89
Qwen-VL-Chat 32.17 $\pm$ 0.24 40.67 $\pm$ 1.03 58.00 $\pm$ 0.71 84.67 $\pm$ 1.03 88.33 $\pm$ 2.05
LLaVA-Next-7B 20.50 $\pm$ 0.00 64.50 $\pm$ 0.82 52.83 $\pm$ 1.25 7.83 $\pm$ 1.65 7.83 $\pm$ 1.55
Emu2-Chat 29.89 $\pm$ 0.00 50.17 $\pm$ 1.44 51.43 $\pm$ 1.52 59.38 $\pm$ 2.03 57.25 $\pm$ 3.06
GPT4V 48.00 56.00 78.00 90.00 86.00
Table 14: Results of different models on Operator Induction dataset (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
OpenFlamingo-9B 5.00 $\pm$ 0.00 2.22 $\pm$ 3.14 1.67 $\pm$ 1.36 2.78 $\pm$ 0.79 7.78 $\pm$ 2.08
IDEFICS-9B 11.67 $\pm$ 0.00 14.44 $\pm$ 0.79 10.56 $\pm$ 2.08 7.78 $\pm$ 2.08 $1$$\pm$ 1.57
IDEFICS-80B 13.33 $\pm$ 0.00 15.00 $\pm$ 2.72 14.67 $\pm$ 2.36 21.67 $\pm$ 1.36 16.11 $\pm$ 2.08
Otter 21.67 $\pm$ 0.00 11.67 $\pm$ 2.36 13.33 $\pm$ 1.36 12.22 $\pm$ 1.57 7.22 $\pm$ 1.57
InternLM-X2 26.11 $\pm$ 3.14 40.00 $\pm$ 10.80 40.00 $\pm$ 4.91 39.44 $\pm$ 7.49 28.89 $\pm$ 19.83
Qwen-VL-Chat 15.00 $\pm$ 0.00 10.00 $\pm$ 1.36 17.22 $\pm$ 3.14 18.89 $\pm$ 1.57 25.00 $\pm$ 2.72
LLaVA-Next-7B 10.56 $\pm$ 1.57 6.11 $\pm$ 1.57 5.56 $\pm$ 2.08 3.33 $\pm$ 2.72 0.00 $\pm$ 0.00
Emu2-Chat 28.56 $\pm$ 1.57 21.67 $\pm$ 5.93 21.11 $\pm$ 1.57 21.67 $\pm$ 0.00 21.11 $\pm$ 5.50
GPT4V 24.00 66.00 84.00 92.00 92.00
Table 15: Results of different models on TextOCR dataset (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
OpenFlamingo-9B 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
IDEFICS-9B 16.50 $\pm$ 0.00 22.50 $\pm$ 1.08 19.83 $\pm$ 0.62 22.83 $\pm$ 1.31 28.00 $\pm$ 1.63
IDEFICS-80B 20.00 $\pm$ 0.00 25.50 $\pm$ 2.18 25.38 $\pm$ 2.78 29.50 $\pm$ 2.89 23.50 $\pm$ 3.47
Otter 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.17 $\pm$ 0.24 0.83 $\pm$ 0.47 0.67 $\pm$ 0.24
InternLM-X2 8.67 $\pm$ 4.01 3.83 $\pm$ 0.62 10.50 $\pm$ 0.71 16.00 $\pm$ 2.48 11.83 $\pm$ 2.95
Qwen-VL-Chat 4.83 $\pm$ 6.84 17.17 $\pm$ 1.43 21.50 $\pm$ 1.08 22.33 $\pm$ 1.31 24.17 $\pm$ 0.24
LLaVA-Next-7B 24.67 $\pm$ 2.25 0.83 $\pm$ 0.24 0.33 $\pm$ 0.24 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Emu2-Chat 25.83 $\pm$ 0.24 23.50 $\pm$ 1.47 31.50 $\pm$ 1.87 36.50 $\pm$ 1.87 29.50 $\pm$ 1.78
GPT4V 39.29 32.14 48.00 50.00 49.00
Table 16: Results of different models on CLEVR dataset (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
OpenFlamingo-9B 0.00 $\pm$ 0.00 17.83 $\pm$ 2.25 17.00 $\pm$ 2.27 18.83 $\pm$ 1.03 16.33 $\pm$ 1.43
IDEFICS-9B 0.00 $\pm$ 0.00 30.33 $\pm$ 2.25 29.50 $\pm$ 1.47 27.67 $\pm$ 2.05 27.17 $\pm$ 2.87
IDEFICS-80B 0.00 $\pm$ 0.00 31.16 $\pm$ 2.10 30.82 $\pm$ 1.59 31.50 $\pm$ 1.00 32.43 $\pm$ 3.62
Otter 0.00 $\pm$ 0.00 5.42 $\pm$ 1.06 8.33 $\pm$ 2.24 8.17 $\pm$ 1.44 0.17 $\pm$ 0.24
InternLM-X2 1.83 $\pm$ 0.24 26.00 $\pm$ 1.63 24.67 $\pm$ 5.25 20.00 $\pm$ 2.94 22.83 $\pm$ 0.85
Qwen-VL-Chat 0.00 $\pm$ 0.00 29.83 $\pm$ 4.55 25.33 $\pm$ 3.47 26.83 $\pm$ 3.06 30.17 $\pm$ 2.95
LLaVA-Next-7B 0.00 $\pm$ 0.00 25.17 $\pm$ 6.64 24.83 $\pm$ 4.90 17.83 $\pm$ 4.59 0.17 $\pm$ 0.24
Emu2-Chat 5.33 $\pm$ 0.24 11.83 $\pm$ 2.72 14.00 $\pm$ 3.49 14.83 $\pm$ 1.89 17.67 $\pm$ 1.03
GPT4V 6.00 30.00 38.00 42.00 32.00
Table 17: Results of different models on Interleaved Operator induction (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
OpenFlamingo-9B 0.00 $\pm$ 0.00 5.56 $\pm$ 1.57 3.89 $\pm$ 2.83 2.78 $\pm$ 0.79 8.89 $\pm$ 3.42
IDEFICS-9B 15.00 $\pm$ 0.00 5.56 $\pm$ 2.08 6.11 $\pm$ 0.79 6.11 $\pm$ 1.57 5.00 $\pm$ 2.36
IDEFICS-80B 25.00 $\pm$ 0.00 36.67 $\pm$ 1.21 31.67 $\pm$ 2.46 28.33 $\pm$ 3.13 20.00 $\pm$ 2.77
Otter 8.33 $\pm$ 0.00 7.78 $\pm$ 1.57 9.44 $\pm$ 3.14 7.22 $\pm$ 2.83 5.56 $\pm$ 2.83
InternLM-X2 28.33 $\pm$ 0.00 10.56 $\pm$ 2.83 9.44 $\pm$ 2.83 $1$$\pm$ 3.93 4.44 $\pm$ 2.83
Qwen-VL-Chat 16.67 $\pm$ 0.00 9.44 $\pm$ 0.79 8.33 $\pm$ 1.36 8.89 $\pm$ 2.83 5.56 $\pm$ 0.79
LLaVA-Next-7B 13.89 $\pm$ 1.57 7.22 $\pm$ 2.83 6.11 $\pm$ 3.14 5.00 $\pm$ 0.00 5.00 $\pm$ 2.72
Emu2-Chat 26.67 $\pm$ 0.00 18.33 $\pm$ 2.72 20.56 $\pm$ 3.42 10.00 $\pm$ 0.00 7.62 $\pm$ 1.83
GPT4V 36.00 58.00 72.00 74.00 70.00
Table 18: Results of different models on Matching MiniImageNet dataset (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 5-Shot
OpenFlamingo-9B 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00
IDEFICS-9B 50.00 $\pm$ 0.00 50.50 $\pm$ 0.35 50.83 $\pm$ 0.85 50.00 $\pm$ 0.20 49.92 $\pm$ 0.12
IDEFICS-80B 61.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00
Otter 48.75 $\pm$ 0.00 50.58 $\pm$ 0.31 50.92 $\pm$ 0.42 50.42 $\pm$ 0.12 49.83 $\pm$ 0.31
InternLM-XComposer2 63.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.08 $\pm$ 1.65 50.00 $\pm$ 0.00
Qwen-VL-Chat 50.50 $\pm$ 0.50 57.32 $\pm$ 1.82 55.50 $\pm$ 1.50 56.43 $\pm$ 1.17 52.82 $\pm$ 0.49
LLaVA-Next-7B 63.00 $\pm$ 0.00 50.00 $\pm$ 0.00 49.75 $\pm$ 0.00 50.00 $\pm$ 0.00 49.75 $\pm$ 0.00
Emu2-Chat 62.15 $\pm$ 3.28 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00
GPT4V 52.00 76.00 82.00 81.00 82.00
Table 19: Results of different models on CobSAT: Total accuracies (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
GILL 2.67 $\pm$ 0.24 12.33 $\pm$ 1.31 9.33 $\pm$ 0.24 11.50 $\pm$ 1.47 8.00 $\pm$ 1.63
SEED-LLaMA-8B 0.50 $\pm$ 0.41 15.83 $\pm$ 1.65 21.83 $\pm$ 1.65 27.83 $\pm$ 2.36 33.67 $\pm$ 2.32
SEED-LLaMA-14B 5.50 $\pm$ 0.71 26.83 $\pm$ 1.65 $3$$\pm$ 3.32 40.83 $\pm$ 1.65 43.83 $\pm$ 2.87
Emu1-Gen 0.33 $\pm$ 0.47 4.83 $\pm$ 0.47 6.17 $\pm$ 2.72 8.67 $\pm$ 1.18 9.67 $\pm$ 0.24
Emu2-Gen 8.67 $\pm$ 0.62 23.00 $\pm$ 3.24 28.67 $\pm$ 2.01 27.33 $\pm$ 2.72 20.83 $\pm$ 0.85
Table 20: Results of different models on CobSAT: Latent accuracies (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
GILL 7.67 $\pm$ 0.24 47.67 $\pm$ 1.43 53.17 $\pm$ 1.65 67.33 $\pm$ 1.03 72.33 $\pm$ 0.85
SEED-LLaMA-8B 8.00 $\pm$ 0.41 38.50 $\pm$ 1.47 44.33 $\pm$ 0.62 49.50 $\pm$ 0.82 56.00 $\pm$ 1.41
SEED-LLaMA-14B 15.17 $\pm$ 0.62 41.50 $\pm$ 1.41 52.17 $\pm$ 2.09 53.67 $\pm$ 1.93 57.17 $\pm$ 1.84
Emu1-Gen 8.00 $\pm$ 0.41 55.50 $\pm$ 2.55 71.00 $\pm$ 0.71 77.00 $\pm$ 0.82 82.00 $\pm$ 0.00
Emu2-Gen 18.00 $\pm$ 1.63 43.83 $\pm$ 4.33 72.33 $\pm$ 1.25 81.50 $\pm$ 0.41 78.33 $\pm$ 1.25
Table 21: Results of different models on CobSAT: Non-latent accuracies (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
GILL 19.33 $\pm$ 0.24 21.33 $\pm$ 1.65 16.17 $\pm$ 1.65 19.83 $\pm$ 2.01 15.83 $\pm$ 2.78
SEED-LLaMA-8B 12.33 $\pm$ 1.03 47.33 $\pm$ 4.03 52.00 $\pm$ 1.87 58.83 $\pm$ 1.93 63.33 $\pm$ 2.01
SEED-LLaMA-14B 82.67 $\pm$ 0.24 76.33 $\pm$ 0.94 75.83 $\pm$ 1.70 78.33 $\pm$ 0.85 80.83 $\pm$ 0.85
Emu1-Gen 26.50 $\pm$ 0.41 11.17 $\pm$ 0.47 13.33 $\pm$ 2.25 16.00 $\pm$ 0.71 17.00 $\pm$ 0.71
Emu2-Gen 62.00 $\pm$ 0.41 49.17 $\pm$ 4.29 42.33 $\pm$ 2.62 35.67 $\pm$ 2.05 29.33 $\pm$ 1.43
Table 22: Results of different models on Text-to-Image Fast Mini-ImageNet (Accuracy %)
Model 0-Shot 1-Shot 2-Shot 4-Shot 5-Shot
GILL 0.00 $\pm$ 0.00 16.00 $\pm$ 2.27 15.17 $\pm$ 2.72 14.83 $\pm$ 0.24 14.33 $\pm$ 2.25
SEED-LLaMA-8B 0.00 $\pm$ 0.00 15.00 $\pm$ 3.27 12.67 $\pm$ 1.18 16.00 $\pm$ 2.12 16.50 $\pm$ 1.87
SEED-LLaMA-14B 0.75 $\pm$ 0.25 17.25 $\pm$ 2.75 16.75 $\pm$ 1.75 21.25 $\pm$ 1.75 21.00 $\pm$ 3.00
Emu1-Gen 0.50 $\pm$ 0.41 31.50 $\pm$ 1.87 22.83 $\pm$ 2.72 25.00 $\pm$ 0.71 23.17 $\pm$ 1.03
Emu2-Gen 0.00 $\pm$ 0.00 24.33 $\pm$ 3.30 30.67 $\pm$ 1.31 37.00 $\pm$ 1.22 34.50 $\pm$ 0.00
Table 23: Results of different models on the Text version of Operator Induction (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
InternLM-XComposer2 15.00 50.00 73.33 75.00 83.33
Qwen-VL-Chat 0.00 45.00 56.67 63.33 71.67
LLaVA-Next-Vicuna-7B 10.00 40.00 53.33 60.00 68.33
Table 24: Results of different models on the text version of Interleaved Operator Induction (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
InternLM-XComposer2 8.33 35.00 36.67 46.67 78.33
Qwen-VL-Chat 0.00 50.00 55.00 61.67 66.67
LLaVA-Next-Vicuna-7B 16.67 41.67 45.00 53.33 70.00
3.3 Additional Analysis
We present the results of the additional analysis in Table 25 to 34.
Table 25: Results of different models on the text version of CLEVR dataset (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
InternLM-XComposer2 0.00 45.00 43.00 42.00 41.00
Qwen-VL-Chat 0.00 49.50 47.50 54.00 53.50
LLaVA-Next-Vicuna-7B 0.00 43.00 38.50 37.50 36.50
Table 26: Results of different models on the text version of Text-to-Image Fast Mini-ImageNet (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 5-Shot
GILL 0.00 18.50 20.00 20.50 18.50
SEED-LLaMA-8B 0.00 16.30 15.20 16.50 14.20
SEED-LLaMA-14B 1.50 23.00 20.00 22.50 15.50
Emu1-Gen 0.50 28.60 29.10 24.20 20.00
Emu2-Gen 0.20 32.40 38.80 40.50 42.10
Table 27: Results of different models on the text version of CobSAT: Total accuracies (%).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
GILL 6.00 13.00 20.50 22.50 23.50
SEED-LLaMA-8B 0.50 14.50 15.50 30.50 32.00
SEED-LLaMA-14B 6.00 13.50 28.00 34.00 40.50
Emu1-Gen 2.50 11.00 19.50 23.50 20.00
Emu2-Gen 7.50 19.50 32.50 46.50 45.00
Table 28: Results of different models on the text version of CobSAT: Latent accuracies (%).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
GILL 6.50 33.00 39.00 37.50 38.00
SEED-LLaMA-8B 4.00 17.50 18.50 35.00 46.00
SEED-LLaMA-14B 6.50 60.00 55.50 60.00 66.00
Emu1-Gen 6.00 24.00 31.50 43.50 42.00
Emu2-Gen 12.00 74.00 86.00 92.50 88.50
Table 29: Results of different models on the text version of CobSAT: Non-latent accuracies (%).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
GILL 86.00 44.50 62.50 67.00 71.50
SEED-LLaMA-8B 21.00 80.00 83.50 80.50 74.50
SEED-LLaMA-14B 90.00 21.50 56.50 63.00 67.50
Emu1-Gen 30.00 33.50 52.00 48.50 45.50
Emu2-Gen 68.50 22.50 37.50 50.50 49.00
Table 30: Comparisons of in-context learning ability with and without instruction-following fine-tuning on Fast Open-Ended MiniImageNet Dataset.
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
IDEFICS-9B 0.00 16.00 47.50 58.00 56.00
IDEFICS-9B-Instruct 0.00 22.00 52.00 53.83 59.17
Qwen-VL 0.00 35.50 79.50 68.00 67.00
Qwen-VL-Chat 0.00 0.50 47.33 58.00 55.17
Table 31: Comparisons of in-context learning ability with and without instruction-following fine-tuning on TextOCR Dataset (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
IDEFICS-9B 3.50 16.50 22.50 25.00 26.00
IDEFICS-9B-Instruct 16.50 22.50 19.83 22.83 28.00
Qwen-VL 0.00 27.00 28.50 30.50 37.00
Qwen-VL-Chat 4.83 17.17 21.50 22.33 24.17
Table 32: Comparisons of in-context learning ability with and without instruction-following fine-tuning on CLEVR Dataset (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
IDEFICS-9B 0.00 19.50 26.00 25.00 29.00
IDEFICS-9B-Instruct 0.00 30.33 29.50 27.67 27.17
Qwen-VL 2.50 18.50 17.50 24.00 26.00
Qwen-VL-Chat 0.00 29.83 25.33 26.83 30.17
Table 33: Comparisons of in-context learning ability with and without instruction-following fine-tuning on Operator Induction Dataset (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
IDEFICS-9B 5.00 16.67 8.33 10.00 3.33
IDEFICS-9B-Instruct 11.67 14.44 10.56 7.78 11.11
Qwen-VL 15.00 26.67 36.67 46.67 56.67
Qwen-VL-Chat 15.00 10.00 17.22 18.89 25.00
Table 34: Comparisons of in-context learning ability with and without instruction-following fine-tuning on Interleaved Operator Induction Dataset (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
IDEFICS-9B 13.33 8.33 5.00 10.00 3.33
IDEFICS-9B-Instruct 15.00 5.56 6.11 6.11 5.00
Qwen-VL 0.00 13.33 13.33 8.33 11.67
Qwen-VL-Chat 16.67 9.44 8.33 8.89 5.56
3.4 Supplementary Results
Below we present the raw results of the figures in the appendix.
3.4.1 Scaling to Many Shots
Table 35 to 38
3.4.2 Chain-of-Thought Prompting
Table 39 to 43
3.4.3 Repeating Support Set
Table 44 to 46.
3.4.4 Different Levels of Task Descriptions
Table 47 to 50.
Table 35: Results of many shots on CLEVR dataset.
Model 16-Shot 32-Shot 64-Shot
OpenFlamingo-9B 22.00 $\pm$ 1.47 25.33 $\pm$ 1.65 25.67 $\pm$ 2.39
IDEFICS-9B-Instruct 28.17 $\pm$ 2.66 29.00 $\pm$ 1.08 30.50 $\pm$ 1.78
InternLM-X2 14.67 $\pm$ 1.70 15.50 $\pm$ 1.08 16.33 $\pm$ 1.03
Table 36: Results of many shots on Operator Induction dataset.
Model 16-Shot 32-Shot 64-Shot
OpenFlamingo-9B 13.33 $\pm$ 3.60 8.89 $\pm$ 1.57 11.67 $\pm$ 1.36
IDEFICS-9B-Instruct 5.00 $\pm$ 3.60 7.78 $\pm$ 1.57 5.00 $\pm$ 1.36
InternLM-X2 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Table 37: Results of many shots on Interleaved Operator Induction dataset.
Model 16-Shot 32-Shot 64-Shot
OpenFlamingo-9B 8.89 $\pm$ 3.42 8.33 $\pm$ 3.60 11.67 $\pm$ 3.60
IDEFICS-9B-Instruct 8.89 $\pm$ 2.08 7.78 $\pm$ 2.08 7.78 $\pm$ 2.83
InternLM-X2 3.89 $\pm$ 0.79 5.00 $\pm$ 1.36 5.00 $\pm$ 1.36
Table 38: Results of many shots on TextOCR dataset.
Model 16-Shot 32-Shot 64-Shot
OpenFlamingo-9B 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
IDEFICS-9B-Instruct 29.00 $\pm$ 1.22 33.17 $\pm$ 0.85 33.50 $\pm$ 1.47
InternLM-X2 3.11 $\pm$ 0.58 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Table 39: Results with Chain-of-Thought prompting on Operator Induction dataset.
Model 0-shot 1-shot 2-Shot 4-Shot 8-Shot
Qwen-VL-Chat 13.67 21.00 15.33 9.33 16.67
InternLM-X2 28.33 32.00 35.67 30.00 5.00
Table 40: Results with Chain-of-Thought prompting on Interleaved Operator Induction dataset.
Model 0-shot 1-shot 2-Shot 4-Shot 8-Shot
Qwen-VL-Chat 3.33 10.00 10.00 8.33 8.33
InternLM-X2 18.33 5.00 5.00 10.00 8.33
Table 41: Results with Chain-of-Thought prompting on TextOCR dataset.
Model 0-shot 1-shot 2-Shot 4-Shot 8-Shot
Qwen-VL-Chat 7.00 28.50 27.50 30.50 22.50
InternLM-X2 12.00 1.50 0.50 2.00 0.50
Table 42: Results with Chain-of-Thought prompting on CLEVR dataset.
Model 0-shot 1-shot 2-Shot 4-Shot 8-Shot
Qwen-VL-Chat 0.50 10.50 21.50 18.00 26.00
InternLM-X2 4.00 21.50 20.00 26.00 27.00
Table 43: Results with Chain-of-Thought prompting on Matching Mini-ImageNet dataset.
Model 0-shot 1-shot 2-Shot 4-Shot 5-Shot
Qwen-VL-Chat 56.00 56.75 51.25 56.75 53.00
InternLM-X2 58.25 51.50 53.00 50.00 48.50
Table 44: Result of Qwen-VL-Chat on Fast Open-Ended MiniImageNet dataset with repeated in-context examples.
Model 1-shot 2-Shot 4-Shot 5-Shot
No Repeat 0.50 47.33 58.00 55.17
Repeat x2 41.00 62.50 54.50 56.50
Repeat x3 62.50 55.50 61.00 62.00
Repeat x4 60.00 56.50 60.00 58.50
Table 45: Result of Qwen-VL-Chat on Operator Induction dataset with repeated in-context examples.
Model 1-shot 2-Shot 4-Shot 8-Shot
No repeat 10.00 17.22 18.89 25.00
Repeat x2 5.00 15.00 23.33 25.00
Repeat x3 11.67 15.00 20.00 26.67
Repeat x4 13.33 18.33 21.67 18.33
Table 46: Result of Qwen-VL-Chat on CLEVR dataset with repeated in-context examples.
Model 1-shot 2-Shot 4-Shot 8-Shot
No Repeat 29.83 25.33 26.83 30.17
Repeat x2 25.50 30.50 27.50 28.50
Repeat x3 22.50 32.50 26.50 32.00
Repeat x4 19.50 31.50 23.00 27.50
Table 47: Results of different models on Fast Open-Ended MiniImageNet dataset (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 5-Shot
Qwen-VL-Chat – Detailed 0.00 $\pm$ 0.00 0.50 $\pm$ 0.41 47.33 $\pm$ 2.49 58.00 $\pm$ 2.83 55.17 $\pm$ 2.25
Qwen-VL-Chat – Concise 0.00 $\pm$ 0.00 0.83 $\pm$ 0.62 48.00 $\pm$ 2.45 59.00 $\pm$ 0.41 52.50 $\pm$ 2.68
Qwen-VL-Chat – None 0.00 $\pm$ 0.00 6.33 $\pm$ 0.47 56.17 $\pm$ 1.65 57.67 $\pm$ 0.85 53.83 $\pm$ 2.78
LLaVA-Next-7B – Detailed 0.00 $\pm$ 0.00 22.17 $\pm$ 4.03 33.67 $\pm$ 2.25 0.00 $\pm$ 0.00 0.33 $\pm$ 0.24
LLaVA-Next-7B – Concise 0.00 $\pm$ 0.00 24.00 $\pm$ 0.71 34.50 $\pm$ 2.68 0.00 $\pm$ 0.00 0.33 $\pm$ 0.24
LLaVA-Next-7B – None 0.00 $\pm$ 0.00 16.67 $\pm$ 2.01 32.00 $\pm$ 2.55 0.33 $\pm$ 0.24 0.17 $\pm$ 0.24
OpenFlamingo-9B – Detailed 0.00 $\pm$ 0.00 39.50 $\pm$ 1.22 58.17 $\pm$ 3.57 51.17 $\pm$ 0.85 54.50 $\pm$ 5.66
OpenFlamingo-9B – Concise 0.00 $\pm$ 0.00 36.50 $\pm$ 0.41 51.67 $\pm$ 2.78 52.17 $\pm$ 0.62 49.33 $\pm$ 1.25
OpenFlamingo-9B – None 0.00 $\pm$ 0.00 38.17 $\pm$ 1.03 52.17 $\pm$ 2.46 49.17 $\pm$ 0.85 49.33 $\pm$ 1.25
InternLM-X2 – Detailed 0.00 $\pm$ 0.00 14.83 $\pm$ 1.03 38.00 $\pm$ 1.78 49.00 $\pm$ 1.78 50.33 $\pm$ 3.86
InternLM-X2 – Concise 0.00 $\pm$ 0.00 19.50 $\pm$ 1.47 40.33 $\pm$ 1.89 48.83 $\pm$ 0.85 49.17 $\pm$ 1.93
InternLM-X2 – None 0.00 $\pm$ 0.00 22.00 $\pm$ 2.04 43.00 $\pm$ 2.16 46.33 $\pm$ 3.06 48.17 $\pm$ 0.62
IDEFICS-9B – Detailed 0.00 $\pm$ 0.00 22.00 $\pm$ 0.41 52.00 $\pm$ 2.94 53.83 $\pm$ 0.94 59.17 $\pm$ 6.20
IDEFICS-9B – Concise 0.00 $\pm$ 0.00 28.50 $\pm$ 1.78 53.83 $\pm$ 4.09 53.83 $\pm$ 0.94 55.67 $\pm$ 2.09
IDEFICS-9B – None 0.00 $\pm$ 0.00 37.17 $\pm$ 4.29 52.17 $\pm$ 4.48 53.17 $\pm$ 1.25 55.50 $\pm$ 1.47
Table 48: Results of different models on CLEVR Count Induction dataset using different levels of task description (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
Qwen-VL-Chat – Detailed 0.00 $\pm$ 0.00 29.83 $\pm$ 4.55 25.33 $\pm$ 3.47 26.83 $\pm$ 3.06 30.17 $\pm$ 2.95
Qwen-VL-Chat – Concise 0.00 $\pm$ 0.00 24.67 $\pm$ 2.32 25.67 $\pm$ 0.85 25.33 $\pm$ 1.65 24.83 $\pm$ 2.32
Qwen-VL-Chat – None 1.00 $\pm$ 0.00 25.17 $\pm$ 2.72 24.33 $\pm$ 1.31 24.83 $\pm$ 1.31 24.67 $\pm$ 2.36
LLaVA-Next-7B – Detailed 0.00 $\pm$ 0.00 25.17 $\pm$ 6.64 24.83 $\pm$ 4.90 17.83 $\pm$ 4.59 0.17 $\pm$ 0.24
LLaVA-Next-7B – Concise 0.00 $\pm$ 0.00 25.00 $\pm$ 3.49 27.00 $\pm$ 3.89 20.00 $\pm$ 2.48 0.00 $\pm$ 0.00
LLaVA-Next-7B – None 0.00 $\pm$ 0.00 15.50 $\pm$ 2.12 23.83 $\pm$ 2.87 12.83 $\pm$ 1.70 0.17 $\pm$ 0.24
OpenFlamingo-9B – Detailed 0.00 $\pm$ 0.00 17.83 $\pm$ 2.25 17.00 $\pm$ 2.27 18.83 $\pm$ 1.03 16.33 $\pm$ 1.43
OpenFlamingo-9B – Concise 0.00 $\pm$ 0.00 15.33 $\pm$ 2.39 19.00 $\pm$ 2.27 20.00 $\pm$ 0.71 18.33 $\pm$ 3.09
OpenFlamingo-9B – None 0.00 $\pm$ 0.00 15.33 $\pm$ 0.94 18.17 $\pm$ 1.03 21.33 $\pm$ 1.89 19.33 $\pm$ 2.78
InternLM-X2 – Detailed 1.83 $\pm$ 0.24 26.00 $\pm$ 1.63 24.67 $\pm$ 5.25 20.00 $\pm$ 2.94 22.83 $\pm$ 0.85
InternLM-X2 – Concise 1.00 $\pm$ 0.00 19.33 $\pm$ 2.25 20.17 $\pm$ 1.31 9.50 $\pm$ 1.41 12.33 $\pm$ 2.32
InternLM-X2 – None 1.50 $\pm$ 0.00 26.67 $\pm$ 2.09 24.67 $\pm$ 2.01 25.17 $\pm$ 1.18 23.17 $\pm$ 2.25
IDEFICS-9B – Detailed 0.00 $\pm$ 0.00 30.33 $\pm$ 2.25 29.50 $\pm$ 1.47 27.67 $\pm$ 2.05 27.17 $\pm$ 2.87
IDEFICS-9B – Concise 1.00 $\pm$ 0.00 30.67 $\pm$ 1.84 31.00 $\pm$ 3.94 26.17 $\pm$ 1.55 26.83 $\pm$ 0.62
IDEFICS-9B – None 0.00 $\pm$ 0.00 30.83 $\pm$ 1.43 31.33 $\pm$ 2.95 28.50 $\pm$ 1.78 28.00 $\pm$ 0.41
Table 49: Results of different models on Operator Induction dataset using different levels of task description (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
Qwen-VL-Chat – Detailed 15.00 $\pm$ 0.00 10.00 $\pm$ 1.36 17.22 $\pm$ 3.14 18.89 $\pm$ 1.57 25.00 $\pm$ 2.72
Qwen-VL-Chat – Concise 15.00 $\pm$ 0.00 7.22 $\pm$ 2.08 15.56 $\pm$ 3.42 17.78 $\pm$ 2.08 27.22 $\pm$ 0.79
Qwen-VL-Chat – None 15.00 $\pm$ 0.00 8.33 $\pm$ 2.36 14.44 $\pm$ 2.83 18.33 $\pm$ 2.72 27.22 $\pm$ 0.79
LLaVA-Next-7B – Detailed 10.56 $\pm$ 1.57 6.11 $\pm$ 1.57 5.56 $\pm$ 2.08 3.33 $\pm$ 2.72 0.00 $\pm$ 0.00
LLaVA-Next-7B – Concise 5.00 $\pm$ 0.00 7.22 $\pm$ 0.79 5.56 $\pm$ 2.08 4.44 $\pm$ 2.08 1.11 $\pm$ 0.79
LLaVA-Next-7B – None 8.33 $\pm$ 0.00 6.11 $\pm$ 0.79 5.56 $\pm$ 1.57 4.44 $\pm$ 1.57 0.56 $\pm$ 0.79
OpenFlamingo-9B – Detailed 5.00 $\pm$ 0.00 2.22 $\pm$ 3.14 1.67 $\pm$ 1.36 2.78 $\pm$ 0.79 7.78 $\pm$ 2.08
OpenFlamingo-9B – Concise 6.67 $\pm$ 0.00 5.00 $\pm$ 3.60 4.44 $\pm$ 3.14 4.44 $\pm$ 1.57 9.44 $\pm$ 1.57
OpenFlamingo-9B – None 6.67 $\pm$ 0.00 5.00 $\pm$ 3.60 3.33 $\pm$ 2.36 4.44 $\pm$ 2.08 11.67 $\pm$ 3.60
InternLM-X2 – Detailed 26.11 $\pm$ 3.14 40.00 $\pm$ 10.80 40.00 $\pm$ 4.91 39.44 $\pm$ 7.49 28.89 $\pm$ 19.83
InternLM-X2 – Concise 18.33 $\pm$ 0.00 29.44 $\pm$ 3.42 22.78 $\pm$ 2.83 18.33 $\pm$ 1.36 16.67 $\pm$ 2.36
InternLM-X2 – None 18.33 $\pm$ 0.00 13.33 $\pm$ 2.36 12.78 $\pm$ 2.83 12.22 $\pm$ 2.08 16.67 $\pm$ 2.72
IDEFICS-9B – Detailed 11.67 $\pm$ 0.00 14.44 $\pm$ 0.79 10.56 $\pm$ 2.08 7.78 $\pm$ 2.08 $1$$\pm$ 1.57
IDEFICS-9B – Concise 15.00 $\pm$ 0.00 13.89 $\pm$ 2.83 12.22 $\pm$ 0.79 8.89 $\pm$ 0.79 8.33 $\pm$ 3.60
IDEFICS-9B – None 15.00 $\pm$ 0.00 17.22 $\pm$ 2.83 10.56 $\pm$ 0.79 10.56 $\pm$ 2.08 7.78 $\pm$ 3.93
Table 50: Results of different models on TextOCR dataset using different levels of task description (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 8-Shot
Qwen-VL-Chat – Detailed 4.83 $\pm$ 6.84 17.17 $\pm$ 1.43 21.50 $\pm$ 1.08 22.33 $\pm$ 1.31 24.17 $\pm$ 0.24
Qwen-VL-Chat – Concise 0.00 $\pm$ 0.00 8.00 $\pm$ 0.82 9.50 $\pm$ 0.41 9.83 $\pm$ 0.62 9.17 $\pm$ 0.24
Qwen-VL-Chat – None 0.00 $\pm$ 0.00 9.67 $\pm$ 0.62 10.33 $\pm$ 0.47 10.67 $\pm$ 0.47 9.33 $\pm$ 0.47
LLaVA-Next-7B – Detailed 24.67 $\pm$ 2.25 0.83 $\pm$ 0.24 0.33 $\pm$ 0.24 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
LLaVA-Next-7B – Concise 8.50 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
LLaVA-Next-7B – None 10.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
OpenFlamingo-9B – Detailed 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
OpenFlamingo-9B – Concise 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
OpenFlamingo-9B – None 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
InternLM-X2 – Detailed 8.67 $\pm$ 4.01 3.83 $\pm$ 0.62 10.50 $\pm$ 0.71 16.00 $\pm$ 2.48 11.83 $\pm$ 2.95
InternLM-X2 – Concise 0.50 $\pm$ 0.00 0.50 $\pm$ 0.41 0.83 $\pm$ 0.47 2.33 $\pm$ 1.03 0.00 $\pm$ 0.00
InternLM-X2 – None 0.50 $\pm$ 0.00 0.50 $\pm$ 0.41 1.33 $\pm$ 0.47 3.67 $\pm$ 2.09 0.00 $\pm$ 0.00
IDEFICS-9B – Detailed 16.50 $\pm$ 0.00 22.50 $\pm$ 1.08 19.83 $\pm$ 0.62 22.83 $\pm$ 1.31 28.00 $\pm$ 1.63
IDEFICS-9B – Concise 3.00 $\pm$ 0.00 2.50 $\pm$ 0.41 5.50 $\pm$ 0.41 5.83 $\pm$ 0.24 6.17 $\pm$ 0.47
IDEFICS-9B – None 4.00 $\pm$ 0.00 2.67 $\pm$ 0.62 5.33 $\pm$ 0.47 6.00 $\pm$ 0.41 6.33 $\pm$ 0.62